# Cocoa: Co-Planning and Co-Execution with AI Agents

### v2

@@ A.7. Can I use Cocoa and/or the baseline?
We are interested in continuing to iterate on Cocoa, so we are not yet planning a public release. However, we may provide some researchers with access to Cocoa to in their day-to-day work as part of a longitudinal study. If you are interested in participating in this study, please reach out. The baseline system is tentatively planned for public release in 2025.
@@ C.1. Project document activity
First, we asked participants to discuss some of the ideas in their project document they shared with us, the origins of those ideas, and the ideal ways they envision AI-powered tools helping them throughout the research process. We then invited participants to a Google Doc or Slides that we converted from their project document for easy collaboration; this conversion was made by copying and pasting content from their project document if it was not already in Docs or Slides. We asked participants to brainstorm 2–3 remaining questions they have about their project that they would be interested in tackling. For each question, participants created a bulleted plan they would undertake to pursue it, which we then used to answer RQ2.
@@ C.2. Probe activity
In the second activity, we presented participants with a WoZ design probe in the form of two side-by-side Google Docs. One Doc, the “planning document,” contained 3 research ideas generated by Perplexity AI (Perplexity, 2024) based on the description of research interests submitted by the participant and 1–3 of their most recent publications and/or preprints. Participants were asked to work in the doc to expand upon an idea (or combine multiple) through a self-defined plan, with the option of invoking the help of an AI assistant through requests prefaced with a “!” command. The study facilitator (the Wizard) entered this command into Perplexity AI and pasted the output to the other Doc, the “AI output document.” While waiting for the response, we asked participants to write a brief plan consisting of 3–6 steps for how they would complete the request themselves. Participants then inspected the assistant’s output and incorporated any useful text into their document. This process repeated until the participant felt like their idea was sufficiently concrete to write a short descriptive paragraph about it, typically after at least 2–3 requests to the AI. This activity helped us answer RQ3.
@@ C.3. Concluding interview
Appendix D Cocoa User Study Participants
@@
Appendix E User Study Self-Evaluation Form Questions
@@
Appendix F Prompts
F.1. Plan generation
@@
F.2. Plan description to agent instruction
@@
F.3. Agent output reformatting for displaying in UI

### v3

@@ A.7. Can I use Cocoa and/or the baseline?
We are continuing to iterate on Cocoa and the baseline, so we are not yet planning a public release. However, we encourage you to check out Ai2’s ScholarQA [^12] https://scholarqa.allen.ai/chat and PaperFinder [^13] https://paperfinder.allen.ai/chat tools, which share some similar features.
@@ Appendix B Formative Study Participants
[…]
Table with the participant ID, gender, age range, country in which they conduct their research, years of experience conducting research, specific research area, general research area, and AI use frequency for each participant from our formative study.
@@ C.1. Project document activity
First, we asked participants to discuss some of the ideas in their project document they shared with us, the origins of those ideas, and the ideal ways they envision AI-powered tools helping them throughout the research process. We then invited participants to a Google Doc or Slides that we converted from their project document for easy collaboration; this conversion was made by copying and pasting content from their project document if it was not already in Docs or Slides. We asked participants to brainstorm 2–3 remaining questions they have about their project that they would be interested in tackling. For each question, participants created a bulleted plan they would undertake to pursue it, which we then used to answer F2.
@@ C.2. Probe activity
In the second activity, we presented participants with a WoZ design probe in the form of two side-by-side Google Docs. One Doc, the “planning document,” contained 3 research ideas generated by Perplexity AI (Perplexity, 2024) based on the description of research interests submitted by the participant and 1–3 of their most recent publications and/or preprints. Participants were asked to work in the doc to expand upon an idea (or combine multiple) through a self-defined plan, with the option of invoking the help of an AI assistant through requests prefaced with a “!” command. The study facilitator (the Wizard) entered this command into Perplexity AI and pasted the output to the other Doc, the “AI output document.” While waiting for the response, we asked participants to write a brief plan consisting of 3–6 steps for how they would complete the request themselves. Participants then inspected the assistant’s output and incorporated any useful text into their document. This process repeated until the participant felt like their idea was sufficiently concrete to write a short descriptive paragraph about it, typically after at least 2–3 requests to the AI. This activity helped us answer F3.
@@ C.3. Concluding interview
Appendix D Formative Study Data Analysis
To answer F1, the first author used a hybrid inductive-deductive coding process (Fereday and Muir-Cochrane, 2006) to code participants’ submitted project documents. The first 5 documents were inductively coded to surface common structural elements before the elements were deductively applied to the remaining 4. We iterated on existing elements and added new ones as needed.
For F2, we performed inductive thematic analysis on two documents—participants’ submitted project documents and in-study planning documents. The first author sourced intentions for planning within submitted project documents by locating the snippet of text that initiates a plan. For example, if the plan consists of a bulleted list, the intention is often expressed in the lead-in text that immediately precedes the list. We note that an individual plan item can also signal planning intent if it contains a nested plan. The first author performed thematic analysis by inductively coding the extracted text. The first author then identified and extracted plan steps participants wrote in both documents before inductively coding them.
For F3, the first author performed open coding on the study transcripts before thematically analyzing the coded snippets. To enrich the data, the first author also extracted and inductively coded all requests to the assistant from participants’ planning documents. The codes were discussed and iterated upon with other project team members on a weekly basis over the course of 3 weeks.
Appendix E Lab Study Participants
@@
[…]
Table with the participant ID, gender, age range, country in which they conduct their research, job title, years of experience conducting research, specific research area, general research area, and AI use frequency for each participant from our lab study.
Appendix F Lab Study Recruitment Process and Pilot Studies
We sent out a recruitment form to collect basic demographic details, frequency of AI use in research, and a copy of a project document from an ongoing or completed project document to use during the study. One participant also participated in our formative study. 15 researchers were based in the United States, and 1 was based in Canada. Participants’ research areas ranged from large language models, to ubiquitous computing, to visualization; broadly, they spanned HCI ($n=10$), ML ($n=3$), and NLP ($n=3$). In our participant pool, many users were occasional (6) or frequent [^14] We define “occasional” and “frequent” the same way as we do in our formative study (Table 1). (6) users of AI tools in the research process.
We also recruited 4 additional participants (1 female, 2 male, 1 non-binary) from our institution for pilot studies to test out Cocoa, catch usability issues, and help us refine our procedure. We learned from our pilot studies that the length of generated plans should be around 3 steps (rather than the system’s default of 5–6 steps) to fit within the study’s time constraints.
Appendix G Lab Study Procedural Details
To control for the length of the study, the first author reviewed the participants’ documents prior to the study to identify and highlight two candidate tasks. To make sure the two tasks are valid and comparable, early on in the study, the first author asked participants 1) whether the tasks have been decisively solved by the participant and whether they would still like to explore them in the study, and 2) whether the two tasks are similar in scope, such that one task does not take significantly more resources and effort to explore than the other. If these conditions were not met, the participant was prompted to rewrite one or both tasks until they were. The two tasks were then randomly assigned to Cocoa and the baseline.
In the Cocoa condition, participants first watched a 6-minute video demo of the system, and were then directed to a document where they would get used to using Cocoa’s features on the following sample question: “How can we use AI to better society?” This practice period lasted around 10 minutes. Participants were then asked to spend 25 minutes to make as much progress as possible tackling the task assigned to this condition with Cocoa, thinking aloud as they did so. They were also permitted to also use other tools (academic search engines, other AI tools, social media, etc.) alongside Cocoa, although very few did. After the 25 minute session, participants wrote a brief set of takeaways and next steps from their exploration, and filled out a short evaluation form with 5-point Likert scale questions about their experience (see Appendix H for this form). After submitting the form, they were encouraged to further try out Cocoa on other parts of their project document for another 10–15 minutes.
Appendix H Lab Study Self-Evaluation Form Questions
@@
Appendix I Prompts
I.1. Plan generation
@@
I.2. Plan description to agent instruction
@@
I.3. Agent output reformatting for displaying in UI
@@
Appendix J Deployment Study General Guidelines
  1. (1)
    Try to spend a total of at least 90 minutes with Cocoa during the 7-day use period.
  2. (2)
    It may help to jot down any notable experiences with Cocoa so we can more easily discuss them later.
  3. (3)
    Treat the document as a scratch space to tinker with and explore rough ideas.
  4. (4)
    Cocoa is just another tool in your toolbox. You can use it alongside whatever you typically use in your workflows!
  5. (5)
    If the system crashes at any point, try refreshing the page. The document auto-saves regularly.